OpenStudy (anonymous):

Suzette must commute a total of 34 miles to work. She first drives her car 10 miles to the train station and then rides the train the rest of the way to work. The average speed of the train is 15 mph hfaster than the average speed of Suzette's car. a. Express the time Suzette spends driving and the time she spends on the train as rational expressions. b. It takes Suzette a total of 1 hour to get to work. Write and solve an equation to find the average speed of Suzette's car. Does Suzette spend a greater amount of time driving or on the train? Explain.

OpenStudy (amistre64):

34 - 10 = 25... train goes 15 faster than car... thinking....

OpenStudy (amistre64):

10(car speed) + 24(train train speed) = something....

OpenStudy (amistre64):

train speed = car spped + 15

OpenStudy (amistre64):

10c + 24(c+15) = something